Garland put 2b + 3 dollars in the bank in the first week. The following week he doubled the first week's savings and put that in the bank. The next week he doubled what was in the bank and put that amount in the bank. If he now has $477 in the bank, how much did he put in the bank the first week.
I have got week DISABLED_event_one= 2b +3
I have got week two = 2 (2b+3)
I am stuck with the third week. Where am I going wrong? I thought maybe 8b+12 = 477 but that does not come out right.
Please help!
Eric

Answer by rothauserc(4718)

week DISABLED_event_one= 2b +3
week two = 2(2b+3) = 4b+6
week three = 4(2b+3) = 8b+12, therefore
477 = 2b+3 +4b+6 +8b+12
477 = 14b +21
14b = 456
b = 32.57
week one he put in 2*32.57 + 3 = $68.14
check this
68.14+(2×68.14)+(4×68.14) = $476.98 is approx $477